OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6. I got 12 and found out it is 4,How?

OpenStudy (debbieg):

It says "find the value of x if g(x) = 6",which means to set g(x)=6 and solve for x. E.g.: 6=3(x - 2) now find x. What YOU did was to let x=6, and then found g(x), e.g., g(6). g(6) is different from "what is x when g(x)=6".

OpenStudy (debbieg):

Remember that the notation g(x) is just function notation for the equation's y-coordinate. It's as if you had the equation: y=3(x - 2) and you were asked: "If y=6, what is x?"

OpenStudy (anonymous):

I get it now, Thank you